OpenStudy (anonymous):

I have a question we are doing quadratic equation and I have an example which I do not think it can be an quadratic equation. x^5+3x-7=0

OpenStudy (anonymous):

Because I believe if you factor the equation it becomes x^5-4

OpenStudy (anonymous):

qudratic form : x^2 + x + C x^5 is def not quadratic

OpenStudy (anonymous):

so in other words you can't solve this equation?

OpenStudy (anonymous):

This x^5-3x+7=0 is in standard form?

OpenStudy (anonymous):

that is not in standard form for quadratic,

OpenStudy (toxicsugar22):

hey can u help me

OpenStudy (anonymous):

suzi20 I have a question for you but if you would take 5^2+3x+7=0 this is a quadratic equation correct?

OpenStudy (anonymous):

5^2+3x+7=0 yea this is quadratic

OpenStudy (anonymous):

the answer would be x=-3+i \[\sqrt{131}/10\]

OpenStudy (anonymous):

and -3-i sqrt 131/10

OpenStudy (anonymous):

suzi20 is that correct what I just posted for the answer

OpenStudy (anonymous):

sorry, i miss read, 5^2+3x+7=0 is not quadratic because 5^2 = 25 and this is constanta

OpenStudy (anonymous):

3x + 32 = 0 x = -32/3

OpenStudy (anonymous):

how would you create a quadratic equation that was not constata?

OpenStudy (anonymous):

5x^+3x-7=0 would this be a quadratic equation since the last sign was changed to minus instead of addition

OpenStudy (anonymous):

suzi20 or would it not make a difference would it still be constata?

OpenStudy (anonymous):

5x^2 + 3x - 7 is quadratic since u have variables x^2 and x 5^2 + 3x - 7 is linear because u just have 1 variable which is x

OpenStudy (anonymous):

then the answer would be x=-3+sqrt149/10, -3-sqrt149/10

OpenStudy (anonymous):

5x^2 + 3x - 7 a = 5 b= 3 c=-7 D = b^2 - 4ac = 9 + 140 = 149 x = -b +/- sqrt (D) over 2a = -3 +/- sqrt (149) over 10 x[1] = -3 + sqrt(149) over 10 x[2] = -3 - sqrt(149) over 10

OpenStudy (anonymous):

Thank you suzi20 you was a big help!